# The Nonlinear Filter Model of Stream Cipher Redivivus

### v3

@@ 4 Concrete Proposals
For any $(e,d)$ in the FAA-profile of $({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m}$, the Berlekamp-Massey step in a fast algebraic attack on the filter generator model has time complexity $\mathcal{O}(ED\log D)$, where $E=\sum_{i=0}^{e}{L\choose i}$ and $D=\sum_{i=0}^{d}{L\choose i}$ (see [31] and Section 3.1.5 of [14]). This complexity dominates the overall time complexity of a fast algebraic attack and ignoring the logarithm term, and we take $T$ to be equal to $ED$. The maximum value of $e$ is one less than the algebraic immunity. The number $N$ of keystream bits required is about $2E$. From Proposition 4, we again assume that ${\sf AI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1})=\lceil(2m+1)/4\rceil$. Recall that for any Boolean function, its fast algebraic immunity is at least one more than its algebraic immunity, and so ${\sf FAI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1}))\geq 1+{\sf AI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1})=1+\lceil(2m+1)/4\rceil$. For the concrete analysis, we use the lower bound $1+\lceil m/2\rceil$ as the actual value of ${\sf FAI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1}))$ Define
$\gamma$ $=$ $\min_{\stackrel{{\scriptstyle 1\leq e\leq\lceil m/2\rceil\rfloor,}}{{d=a-e}}}\left(\sum_{i=0}^{e}{L\choose i}\right)\left(\sum_{i=0}^{d}{L\choose i}\right).$ (14)
@@
Concrete choices of $L$ and $m$.
Given $\kappa$, we obtained representative values of $L$ and $m$. The procedure we followed is to choose the value of $L$ to be the first prime number greater than $2\kappa$, and then for the chosen value of $L$, choose $m$ to be the least integer such that $\beta,\gamma>2^{\kappa}$. This ensures security against (fast) algebraic attacks considered above. The number of variables of the filtering function is $n=2m+1$. For each of the obtained values of $L$ and $n$, we checked whether (9) holds. In each case we found that (9) indeed holds. So the obtained values of $L$ and $m$ ensure security against state guessing attacks. Next, using $\kappa$ and the corresponding values of $L$ and $m$, we computed the maximum value of $B$ such that $B\leq\kappa$, $\alpha>2^{B}$, and (10), (11) and (12) hold. In each of the cases that we considered, it turns out that this maximum value of $B$ is in fact $\kappa$. So for the chosen values of $L$ and $m$, the corresponding $\mathcal{S}(L,m)$ ensures $\kappa$-bit security against the above analysed correlation attacks even when the adversary has access to $2^{\kappa}$ keystream bits. We note, however, that generating $2^{\kappa}$ keystream bits from a single key and IV pair is meaningless from a practical point of view. Instead, we set $B=64$, i.e. from a single key and IV pair at most $2^{64}$ keystream bits are to be generated. Table 1 shows the values of $\kappa$, $L$, $m$ and other parameters of the filtering function. We note the following points regarding the entries in Table 1.
@@
    For each entry in the table, our calculation shows that $\beta\gg\gamma$. For example, for $\mathcal{S}(257,59)$, we obtained $\beta=2^{364.38}$ while $\gamma=2^{130.12}$. This is not surprising, since the fast algebraic attack is expected to be much more faster than the basic algebraic attack.
@@ 5 Efficiency of Computing ${\sf MM}_{n}$
In Section 4, we proposed using $f_{2m+1}$ (which is either ${\sf MM}_{2m+1}$ or $1\oplus{\sf MM}_{2m+1}$) as the filtering function in the nonlinear filter model. Further, Table 1 provides specific suggestions of values of $m$ for achieving different security levels. In this section, we consider the complexity of implementing ${\sf MM}_{2m+1}$. Since ${\sf MM}_{2m+1}$ is constructed from ${\sf MM}_{2m}$ using one XOR gate, we need to consider the complexity of implementing ${\sf MM}_{2m}$.
@@
A half-adder takes two input bits and outputs two bits which represent sum of the two input bits. A full adder takes three input bits and outputs two bits which represent the sum of the three input bits. We estimate the numbers of half and full adders that are required for computing the weight of an $m$-bit string $\mathbf{x}$. The algorithm for computing weight that we use is from [7]. If $m=1$, then no adders are required, if $m=2$, a half adder computes the weight, and if $m=3$, a full adder computes the weight. For $m>3$, write $m=m_{1}+m_{2}+1$, where $m_{1}+1$ is the highest power of two that is at most $m$. The algorithm computes the weight of the first $m_{1}$ bits of $\mathbf{x}$, the weight of the next $m_{2}$ bits of $\mathbf{x}$, and then adds these two weights together together with the last bit of $\mathbf{x}$.
@@
Next we consider concrete estimates. For such estimates, we ignore the at most $\lceil\log_{2}m\rceil$ AND and OR gates required for computing the threshold function from the weight, and also the gates required to compute the next bit of the LFSR. Similar assumptions were made in [1] to obtain the gate count estimate for Grain-128a. To obtain concrete estimates, it is convenient to convert the various gate counts into a single unit. Previous works [1, 9] have taken a single NAND gate as the basic unit and translated other gates in terms of this unit. A half-adder can be implemented using 5 NAND gates, while a full adder can be implemented using 9 NAND gates. In [1, 9], a XOR gate was taken to be 2.5 units and an AND gate was taken to be 1.5 units. Between the papers [1] and [9] there is a difference in the number of units required for a flip-flop:  [1] takes a flip-flop to be 8 units, while [9] takes a flip-flop to be 12 units. In Table 3, we provide estimates of the circuit sizes of $\mathcal{S}(L,m)$ for the values of $(L,m)$ in Table 1. These estimates consider a flip-flop to be 8 units.
$\mathcal{S}(163,36)$ $\mathcal{S}(257,59)$ $\mathcal{S}(331,70)$ $\mathcal{S}(389,87)$ $\mathcal{S}(449,100)$ $\mathcal{S}(521,114)$
@@
If we consider a flip-flop to be 12 units as done for Trivium [9], then the gate count estimate for $\mathcal{S}(163,36)$ is 2395.5. The gate estimate for Trivium obtained in [9] is 3488. Both Trivium and $\mathcal{S}(163,36)$ are targeted at the 80-bit security level, and $\mathcal{S}(163,36)$ is substantially smaller than Trivium. The lower size of $\mathcal{S}(163,36)$ is due to the smaller state size; $\mathcal{S}(163,36)$ uses a 163-bit state, while Trivium uses a 288-bit state.
@@ 6 Conclusion
We described a construction of balanced Boolean functions which has several provable properties, namely very high nonlinearity, acceptable algebraic resistance, and is efficient to implement in hardware. Using such Boolean functions, we proposed concrete constructions of the nonlinear filter model for stream ciphers targeted at different security levels. Gate count estimates for the stream cipher proposals shows that the circuit sizes compare well with well known ciphers at the 80-bit and the 128-bit security levels, while for higher security levels, we do not know of any stream cipher with lower gate count estimates.

### v5

@@ 4.2 Security Analysis
For any $(e,d)$ in the FAA-profile of $({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m}$, the Berlekamp-Massey step in a fast algebraic attack on the filter generator model has time complexity $\mathcal{O}(ED\log D)$, where $E=\sum_{i=0}^{e}{L\choose i}$ and $D=\sum_{i=0}^{d}{L\choose i}$ (see [31] and Section 3.1.5 of [14]). This complexity dominates the overall time complexity of a fast algebraic attack. Ignoring the logarithm term, we take $T$ to be equal to $ED$. The maximum value of $e$ is one less than the algebraic immunity $a$. The number $N$ of keystream bits required is about $2E$. From Proposition 4, we again assume that $a={\sf AI}(({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m+1})=\lceil m/2\rceil$. Define
$\gamma$ $=$ $\min_{\stackrel{{\scriptstyle 1\leq e\leq a-1,}}{{d=a-e}}}\left(\sum_{i=0}^{e}{L\choose i}\right)\left(\sum_{i=0}^{d}{L\choose i}\right).$ (25)
@@
4.3 Concrete Choices of $L$, $m$ and ${\sf pos}$
Given $\kappa$, we obtained representative values of $L$ and $m$. The procedure we followed to obtain $L$ and $m$ is to choose the value of $L$ to be the first prime number greater than $2\kappa$, and then for the chosen value of $L$, choose $m$ to be the least integer such that $\beta,\gamma>2^{\kappa}$. This ensures security against (fast) algebraic attacks considered above. Next, using $\kappa$ and the corresponding values of $L$ and $m$, we computed the maximum value of $B$ such that $B\leq\kappa$, $\alpha>2^{B}$, and (21), (22) and (23) hold. In each of the cases that we considered, it turns out that this maximum value of $B$ is in fact $\kappa$. So for the chosen values of $L$ and $m$, the corresponding $\mathcal{S}(L,m)$ ensures $\kappa$-bit security against the above analysed correlation attacks even when the adversary has access to $2^{\kappa}$ keystream bits. We note, however, that generating $2^{\kappa}$ keystream bits from a single key and IV pair is meaningless from a practical point of view. Instead, we set $B=64$, i.e. from a single key and IV pair at most $2^{64}$ keystream bits are to be generated. Table 1 shows the values of $\kappa$, $L$, $m$ and other parameters of the filtering function. We note the following points regarding the entries in Table 1.
@@ 4.3 Concrete Choices of $L$, $m$ and ${\sf pos}$
    For each entry in the table, our calculation shows that $\beta\gg\gamma$. For example, for $\mathcal{S}(257,59)$, we obtained $\beta=2^{364.38}$ while $\gamma=2^{130.12}$. This is not surprising, since the fast algebraic attack is known to be much more faster than the basic algebraic attack.
@@
Concrete choices of the tap positions.
Given the values of $L$ and $m$ (and also $\kappa$), we applied the procedure for choosing the tap positions encoded by ${\sf pos}$ as outlined in Section 4.1. The leftmost $\kappa$ bits of ${\sf pos}$ is the string ${\sf posX}$. The next $\kappa-1$ bits of ${\sf pos}$ (i.e. the positions ${\sf pos}[L-\kappa-1,\ldots,L-2\kappa+1]$) encode the tap positions for the variables in $\mathbf{Y}$. We define ${\sf posY}$ to be the $\kappa$-bit string formed by appending a 0 to the $(\kappa-1)$-bit segment ${\sf pos}[L-\kappa-1,\ldots,L-2\kappa+1]$ of ${\sf pos}$. So both ${\sf posX}$ and ${\sf posY}$ are $\kappa$-bit strings and are uniquely defined from ${\sf pos}$. Further, given ${\sf posX}$ and ${\sf posY}$ it is possible to uniquely construct ${\sf pos}$ (by concatenating ${\sf posX}$ and ${\sf posY}$, appending $L-2m$ zeros, and setting the bit in position $L-2\kappa$ to be 1, to encode the tap position for $W$). So providing ${\sf posX}$ and ${\sf posY}$ is equivalent to providing ${\sf pos}$. In Table 3 we provide the values of ${\sf posX}$ and ${\sf posY}$ encoded as hexadecimal strings. For example, the entry for ${\sf posX}$ corresponding to $\mathcal{S}(128,59)$ is be352$\cdots$ which encodes the binary string 10111110001101010010$\cdots$. In Table 4 we provide the corresponding values of $\nu$ and $\delta$. Recall that $\nu$ is obtained from ${\sf posX}$ as in (17) and $\delta$ is obtained from ${\sf pos}$ as in (18).
The number of variables of the filtering function is $n=2m+1$. For each of the obtained values of $L$ and $n$, and the corresponding value of $\delta$ we checked whether (20) holds. In each case we found that (20) indeed holds. So the stream cipher proposals ensure security against state guessing attacks.
$\kappa$ ${\sf posX}$ and ${\sf posY}$
$\mathcal{S}(80,37)$ d569a664f500763506c3
ff0149d4c640e9846cf2
$\mathcal{S}(128,59)$ be352d9ca349432b80b38ac54e5164c9
d2ece08cbb5566d608a69b19e4a91418
$\mathcal{S}(160,71)$ ea4308e1229305d185450cfa26b0dcac68c4ab7d
1dbb5a438e7e55904cc04406bf0670ad728462b0
$\mathcal{S}(192,87)$ a0265ea181b73a460fb50d8482590e584d15869de343957e
c6b218be600d6183c074d00fde24e1c308ebb06cebab0f84
$\mathcal{S}(224,101)$ e9507d49d4f4609a710d8d291eb466430af5668b03ec424c18417d86
d288451f8f0554a46615f4448afa34aab8673d0647044afcd4682ec4
$\mathcal{S}(256,115)$ c1ec835120741f290154b122618c625f0a9e77c5172cac84ae564390b2e91fda
5865fda7830eca37d0c2045994e9c83b1c55e13f1966c220809bc019d37f0054
Table 3: The strings ${\sf posX}$ and ${\sf posY}$ corresponding to the values of $L$ and $m$ for $\kappa$ which are shown in Table 1. The first row corresponds to ${\sf posX}$ and the second row corresponds to ${\sf posY}$.
$\mathcal{S}(L,m)$ $\nu$ $\delta$
$\mathcal{S}(80,37)$ 16 36
$\mathcal{S}(128,59)$ 26 57
$\mathcal{S}(160,71)$ 30 69
$\mathcal{S}(192,87)$ 39 86
$\mathcal{S}(224,101)$ 45 96
$\mathcal{S}(256,115)$ 51 112
Table 4: Values of $\nu$ and $\delta$ corresponding to the strings ${\sf posX}$ and ${\sf pos}$ (built from ${\sf posX}$ and ${\sf posY}$) shown in Table 3.
4.4 Differential Attacks
Choosing the filtering function to protect against certain known classes of attacks does not however, protect against all possible attacks. There are attacks which can succeed even if the filtering function is properly chosen. One such attack on a previous version of the proposal was described in [6]. To describe the attack and the design modification to resist it we need to mention the aspects of the design which were different in the previous proposal. There are three such aspects.
  1. 1.
    In the definition of ${\sf MM}_{2m}$, in the earlier version we had chosen $\pi$ to be the identity permutation, whereas in the present version we choose $\pi$ to be the bit reversal permutation.
  2. 2.
    In the previous version, the tap positions were chosen as follows. From the state $(s_{L-1},\ldots,s_{0})$ of the LFSR, the value of $W$ in $f_{2m+1}(W,\mathbf{X},\mathbf{Y})$ was chosen to be the state bit $s_{L-\kappa+m}$, the values of the variables in $\mathbf{X}$ were chosen to be the state bits $s_{L-\kappa+m-1},\ldots,s_{L-\kappa}$, and the values of the variables in $\mathbf{Y}$ were chosen to be the state bits $s_{L-2\kappa+m-1},\ldots,s_{L-2\kappa}$.
  3. 3.
    In the previous version, during the initialisation phase, the feedback from the filtering function was fed back into the state by XORing it with the next bit of the LFSR, i.e. it was fed back to only one position of the state.
Suppose the state after the initialisation phase is $(s_{L-1},\ldots,s_{0})$, and for $t\geq 1$, let $s_{t}={\sf nb}(s_{t-1},\ldots,s_{t-L})$. Let the keystream bit generated from the state $(s_{t},s_{t-1},\ldots,s_{t-L+1})$ be $w_{t}$. Using the definition of $\pi$ to be the identity permutation and the tap positions as mentioned above, we have
$w_{t}$ $=$ $s_{t-\kappa+m}\oplus\langle\pi(s_{t-\kappa+m-1},\ldots,s_{t-\kappa+1},s_{t-\kappa}),(s_{t-2\kappa+m-1},\ldots,s_{t-2\kappa+1},s_{t-2\kappa})\rangle$
$\quad\oplus{\sf Maj}_{m}(s_{t-\kappa+m-1},\ldots,s_{t-\kappa+1},s_{t-\kappa}),$
So
$w_{t}\oplus w_{t+1}$ $=$ $s_{t-\kappa+m-1}s_{t-2\kappa+m-1}\oplus s_{t+1-\kappa}s_{t+1-2\kappa}$
$\oplus{\sf Maj}_{m}(s_{t-\kappa+m-1},\ldots,s_{t-\kappa+1},s_{t-\kappa})\oplus{\sf Maj}_{m}(s_{t-\kappa+m},\ldots,s_{t-\kappa+2},s_{t-\kappa+1}).$
In other words, due to the choice of $\pi$ as the identity permutation, the selection of tap positions as consecutive positions of the state, and the fact that two successive keystream bits are obtained from a single shift of the LFSR sequence, $m-1$ of the quadratic terms in the inner product cancelled out when $w_{t}$ and $w_{t+1}$ are XORed together. Further, the inputs to the two calls to ${\sf Maj}$ have an overlap of size $m-1$. These features were observed in [6] and exploited to mount a differential attack. The idea of the differential attack is to introduce a difference in the top most bit position of the IV. During the initialisation phase, this difference travels unchanged for $\kappa-m$ steps and then produces a difference in the input to the filtering function. The difference in the output of the filtering function is fed back into the leftmost bit of the LFSR. As the LFSR is further shifted, this difference then travels without any further modification creating a high probability truncated differential. Combined with the simplified form of $w_{t}\oplus w_{t+1}$, this leads to an efficient key recovery attack.
In the present version, we have chosen $\pi$ to be the bit reversal permutation, and all the tap positions for $\mathbf{X}$ to occur to the left of all the tap positions for $\mathbf{Y}$. As a result, in the XOR of any number of keystream bits, no quadratic term cancels out (see Proposition 5). So, in particular, the above kind of cancellation of quadratic terms does not arise.
Further, due to our design procedure for the tap positions of corresponding to the variables $X_{1},\ldots,X_{n}$, the overlap of inputs in the calls to ${\sf Maj}$ in the XOR of any two keystream bits is at most $\nu$. From Table 4, we observe that the value of $\nu$ is considerably smaller than $m$. So the complexity of ${\sf Maj}$ is mostly preserved in the XOR of any number of keystream bits.
Further, in the present version, during the initialisation phase we inject the output of the filtering function at multiple positions of the state, the number of positions being controlled by the parameter $\mu$. So any (controlled) difference in any bit position travels at most $\mu$ steps before it is further modified. Since our recommendation is to choose $\mu$ to be about $\sqrt{2\kappa}$, in the full initialisation phase consisting of $2\kappa$ iterations, a difference will be updated about $\sqrt{2\kappa}$ times. This makes it difficult to control a difference through all the initialisation rounds.
Due to the above modifications, the attack in [6] on the previous proposal does not apply to the modified proposal. In fact, the modifications though inexpensive, substantially improve the differential properties of the keystream and also improve the “robustness” of the initialisation phase. We note, however, unlike our analysis of the some of the other attacks, we do not have any proof that our proposal resists all kinds of differential attacks. We welcome further analysis of our proposals, including finding other avenues of attack.
@@ 5 Efficiency of Computing ${\sf MM}_{n}$
In Section 4, we proposed using $f_{2m+1}$ (which is $1\oplus{\sf MM}_{2m+1}$) as the filtering function in the nonlinear filter model. Further, Table 1 provides specific suggestions of values of $m$ for achieving different security levels. In this section, we consider the complexity of implementing ${\sf MM}_{2m+1}$. Since ${\sf MM}_{2m+1}$ is constructed from ${\sf MM}_{2m}$ using one XOR gate, we need to consider the complexity of implementing ${\sf MM}_{2m}$.
@@
A half-adder takes two input bits and outputs two bits which represent sum of the two input bits. A full adder takes three input bits and outputs two bits which represent the sum of the three input bits. We estimate the numbers of half and full adders that are required for computing the weight of an $m$-bit string $\mathbf{x}$. The algorithm for computing weight that we use is from [8]. If $m=1$, then no adders are required, if $m=2$, a half adder computes the weight, and if $m=3$, a full adder computes the weight. For $m>3$, write $m=m_{1}+m_{2}+1$, where $m_{1}+1$ is the highest power of two that is at most $m$. The algorithm computes the weight of the first $m_{1}$ bits of $\mathbf{x}$, the weight of the next $m_{2}$ bits of $\mathbf{x}$, and then adds these two weights together with the last bit of $\mathbf{x}$.
@@
Next we consider concrete estimates. For such estimates, we ignore the at most $\lceil\log_{2}m\rceil$ AND and OR gates required for computing the threshold function from the weight, and the gates required to compute the next bit of the LFSR. Further, we also ignore the $\lceil 2\kappa/\mu\rceil$ XOR gates required for the feedback injection from the filtering function during the initialisation phase. This number is not much. For example, for $\kappa=128$, we have $\mu=\lfloor\sqrt{2\kappa}\rfloor=16$ and so 16 XOR gates are required to feedback the output of the filtering function during the initialisation phase. We obtain concrete estimates for the two major components, the LFSR and the filtering function.
To obtain concrete estimates, it is convenient to convert the various gate counts into a single unit. Previous works [1, 9] have taken a single NAND gate as the basic unit and translated other gates in terms of this unit. A half-adder can be implemented using 5 NAND gates, while a full adder can be implemented using 9 NAND gates. In [1, 9], a XOR gate was taken to be 2.5 units and an AND gate was taken to be 1.5 units. Between the papers [1] and [9] there is a difference in the number of units required for a flip-flop:  [1] takes a flip-flop to be 8 units, while [9] takes a flip-flop to be 12 units. In Table 5, we provide estimates of the circuit sizes of $\mathcal{S}(L,m)$ for the values of $(L,m)$ in Table 1. These estimates consider a flip-flop to be 8 units.
$\mathcal{S}(163,37)$ $\mathcal{S}(257,59)$ $\mathcal{S}(331,70)$ $\mathcal{S}(389,87)$ $\mathcal{S}(449,100)$ $\mathcal{S}(521,114)$
@@
If we consider a flip-flop to be 12 units as done for Trivium [9], then the gate count estimate for $\mathcal{S}(163,37)$ is 2395.5. The gate estimate for Trivium obtained in [9] is 3488. Both Trivium and $\mathcal{S}(163,37)$ are targeted at the 80-bit security level, and $\mathcal{S}(163,37)$ is substantially smaller than Trivium. The lower size of $\mathcal{S}(163,37)$ is due to the smaller state size; $\mathcal{S}(163,37)$ uses a 163-bit state, while Trivium uses a 288-bit state.
@@ 6 Conclusion
We described a construction of balanced Boolean functions which has several provable properties, namely very high nonlinearity, acceptable algebraic resistance, and is efficient to implement in hardware. Using such Boolean functions, we proposed concrete constructions of the nonlinear filter model for stream ciphers targeted at different security levels. Gate count estimates for the stream cipher proposals show that the circuit sizes compare well with famous stream ciphers at the 80-bit and the 128-bit security levels, while for higher security levels, we do not know of any stream cipher with lower gate count estimates.